OpenStudy (jacobbirddog):

Please Help Me... Ill Give You A Medal And A Fan :) What is the slope and y-intercept of this line? y = –6x + 2 A. slope: –6, y-intercept: 2 B. slope: –3, y-intercept: 1 C. slope: 2, y-intercept: –6 D. slope: 6, y-intercept: –2

OpenStudy (jacobbirddog):

Anyone?

OpenStudy (anonymous):

look at this : in all of equations we have : y=mx+d For example : y=5x+8 Slope(m) is the number is before x :5 intercept is d : 8 and y is y .! ;) Now can you answer ?!;)

OpenStudy (jacobbirddog):

umm im not really good at like... algebra...

OpenStudy (anonymous):

Now more help ! Now you SHOULD answer ! y = –6x + 2 y= mx + d m=-6 Whats the d ? m is slope . d is y-intercept . ok ?!:)

OpenStudy (anonymous):

B and C are not answer !

OpenStudy (jacobbirddog):

d @E.ali